OpenStudy (kayla0297):

Exact value of sin(10pi/3)? Can I just convert it to degrees and use the triangle?

OpenStudy (neer2890):

use pi=180 so 1opi/3=10*180/3=600 degrees

OpenStudy (neer2890):

sin(600)=sin(90*6+60)=sin 60

OpenStudy (kayla0297):

Ohhh I completely forgot about doing that, thank you!

OpenStudy (neer2890):

you're welcome...:)